OpenStudy (anonymous):

the population of a country increased in first decade by 20%,30% in the second decade and 45% in the third decade.what is the average rate of increase/decade in the population?

OpenStudy (anonymous):

here geometric mean will be used...

OpenStudy (anonymous):

\[(1.2)(1.3)(1.45) = 2.262\] so increase is 162.2%

OpenStudy (anonymous):

divide by 3 and get average increase i guess

OpenStudy (anonymous):

well satellite u have not been able to help me

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

use the formula sigma logx/n

OpenStudy (anonymous):

and then take the antilog

OpenStudy (anonymous):

if you want the geometric mean it will \[\sqrt[3]{1.2\times 1.3\times 1.45}\]

OpenStudy (anonymous):

hey no this is an alternate formula of g.m in statistics is sigma logx/n

OpenStudy (anonymous):

you want \[\frac{ln(1.2)+\ln(1.3)+\ln(1.45)}{3}\]?

OpenStudy (anonymous):

ya.........

OpenStudy (anonymous):

and what is d logic of using g.m here and not d mean?

OpenStudy (anonymous):

in the book they have also use the compound interst formula alternatively

OpenStudy (anonymous):

ya.......and the wolfram thing is beyond my maths level

OpenStudy (anonymous):

i just plugged in the numbers to the formula you provided. you should be able to see exactly what i typed in

OpenStudy (anonymous):

why did u not take 1.312

hero (hero):

How come you just can't do : (20 + 30 + 45)/3 = 31.6667% average increase?

OpenStudy (anonymous):

because this is the total after the increase. the increase itself is what comes after the one in the decimal

OpenStudy (anonymous):

for example if i want to increase a number by 30% i multiply it by 1.3

OpenStudy (anonymous):

ok ,satellite why didnt we just use the mean?

OpenStudy (anonymous):

gotta run, see you later

OpenStudy (anonymous):

i don't know, you said the book said to use the formula you provided

OpenStudy (anonymous):

bye....i think becoz the population grows exponetially

OpenStudy (anonymous):

i have a similar ques

hero (hero):

tanusingh, have you considered khan academy or patrick JMT?

OpenStudy (anonymous):

ya.......khan academy......but not for dis ques

hero (hero):

http://patrickjmt.com Try that..

hero (hero):

Well, looks like Patrick doesn't have much for Prob and Stats

OpenStudy (anonymous):

lets try hero

OpenStudy (anonymous):

satellite is very gud in maths but he also got stuck

OpenStudy (anonymous):

nah no help

hero (hero):

lol, I am in no position to help you. If Sat can't do it, I don't know who can. He's the smartest guy on here besides James. Now, that's someone you should probably consult.

OpenStudy (anonymous):

herosorry but u r also goood

hero (hero):

I don't even have a stats book. I could probably find one though. Which one are you using?

OpenStudy (anonymous):

ah s.c gupta (indian author

hero (hero):

I don't think I would be able to find that book. I think the solution satellite provided is accurate though.

OpenStudy (anonymous):

statistcal methods by s.c gupta

hero (hero):

I'll try to find it

hero (hero):

There doesn't seem to exist an online version of the book.

OpenStudy (anonymous):

thanks anyway

hero (hero):

I hope you find the help you're looking for. Have you consulted James yet?

hero (hero):

Okay, I'll try to help you but I must ask, "What is your issue with Satellite's solution?".

OpenStudy (anonymous):

why isnt mean used?

OpenStudy (anonymous):

his ans is mathematically right

hero (hero):

He followed the formula you gave him. That's all that's required. I don't believe geometric mean applies to the formula.

hero (hero):

You said it yourself, the formula is an alternative to the geometric mean formula.

OpenStudy (anonymous):

i knw.....

OpenStudy (anonymous):

it is a g.m formula for ungrouped data

hero (hero):

What exactly is it about it that you don't understand? It's just a formula. You plug the numbers in and solve.